# ShapeMoiré: Channel-Wise Shape-Guided Network for Image Demoiréing

### v1

@@ IV-C Ablation Study
The two components we propose, ShapeConv and Shape-Architecture, are independent of each other. This means that adding either one of them to the baseline method alone can enhance the model’s performance. To validate the effectiveness of both components, we conducted detailed ablation studies. The results of these studies are presented in Table III. One can observe that the removal of both ShapeConv and Shape-Architecture leads to degraded model performance. Introducing either the ShapeConv or the Shape-Architecture yields a positive outcome. Combining these two together, i.e., our final ShapeMoiré, achieves the best results, highlighting their complementary effect on the overall model performance.
TABLE III: Ablation study of Shapremoiré on the UHDM dataset.
@@
Additionally, we computed the inference time of both the baseline and our ShapeMoiré method. Specifically, we conducted this experiment on a single NVIDIA A40 GPU with 48G memory and ran both models for three runs. Subsequently, we calculated the average inference time for 5,000 images and presented the milliseconds (ms) per image in Table IV. It is evident that our ShapeMoiré method imposes a very negligible overhead on the baseline model in most cases and even runs faster than the baseline due to clock time oscillation. This results validate the efficiency of the proposed ShapeMoiré method.
TABLE IV: Inference time comparison of baseline and our ShapeMoiré on the UHDM dataset.
@@
[10] ShapeMoiré 1.95 2.07 2.21
IV-D Visualization
In Fig. 4, we present qualitative comparisons between our method and two different baseline approaches, namely WDNet [11] and ESDNet-L [7]. First, we observe that the stronger the baseline method employed, the better the moiré effect removal. This is evident in both Case A and Case B, where employing ESDNet-L as the baseline method yields superior results. Furthermore, ShapeMoiré demonstrates advantages in moiré elimination and image detail restoration regardless of the baseline method’s strength. For instance, in Case A, the sailboat in the top row (b) and its surrounding sky appear to be more faithfully reproduced compared to those in (a). Similarly, in the bottom row of (b) where ESDNet-L is used as the baseline method, the color distribution of the lake surface appears more uniform compared to (a). In Case B, compared to the image in (a), the top row of (b) produced using ShapeMoiré not only eliminates the moiré pattern in the highlighted part of the image but also preserves the texture of the road. Similarly, in the bottom row, where ESDNet-L is the baseline method, the artifacts caused by the incomplete removal of moiré patterns in (b) are smaller compared to (a). All these observations underscore the superiority of our method.
Figure 5: Qualitative results of image demoiréing for real-world cases. The selected samples are drawn from the UHDM dataset. $I^{A}_{moir\acute{e}}$ displays inputs from the dataset’s test samples, while $I^{B}_{moir\acute{e}}$ presenting real photos exhibiting moiré patterns captured using a smartphone. Larger boxes represent amplified views of the corresponding colored small boxes at the borders, offering a clearer perspective when enlarged.
@@ IV-D Visualization
To further validate the practicality of our proposed method, we captured real-world images using smartphones and applied both the baseline and ShapeMoiré methods to these cases, as depicted in Fig. 5. It is important to note that the patterns of moiré are not fixed; for the same original image, different devices may capture different moiré patterns, as illustrated by the two images with moiré, where $I^{A}_{moir\acute{e}}$ is the original input image in the UHDM dataset [7] and $I^{B}_{moir\acute{e}}$ is from our captured smartphone input image. It can be observed that $I^{A}_{moir\acute{e}}$ exhibits more widespread moiré artifacts across the entire region ($patttern^{A}$), while in $I^{B}_{moir\acute{e}}$, the artifacts are concentrated in the sky area at the center of the image ($patttern^{B}$). From the results in Fig. 5, several observations can be made: 1) Our ShapeMoiré consistently outperforms the baseline method, regardless of the two morphologies of moiré, $I^{A}_{moir\acute{e}}$ and $I^{B}_{moir\acute{e}}$; our method exhibits superior moiré removal effects. 2) Our method demonstrates robustness against different moiré patterns, even helping alleviate color distortions. As shown in example $I^{B}_{moir\acute{e}}$ of Fig. 5, due to severe moiré in the blank sky area at the center, noticeable color deviations occur. ShapeMoiré not only excels in removing Moiré patterns but also effectively restores the true colors of the sky.
V Conclusion
@@
In this paper, we extend our previously accepted ShapeConv to the image demoiréing domain. Unlike existing methods, our proposed ShapeMoiré addresses the channel inconsistency problem of moiré patterns using shape features. The key extension in the method is the introduction of the Shape-Architecture, which models the shape information from the image-level and acts as a complement to the patch-level ShapeConv. The experimental results demonstrate its superiority over existing state-of-the-arts on four public popular datasets. Furthermore, we verify that our method can be painlessly integrated into four strong baselines to improve their performance, while introducing negligible computational cost during inference. Additionally, our ShapeMoiré yields significant promise in real-world cases with irregular moiré patterns, making it potentially suitable for edge device application and deployment.
@@ Appendix A Proof for Equation 9
$\mathbb{F}_{c_{out}}=&\sum_{k}^{K_{h}\times K_{w}\times C_{in}}(\mathbb{K}_{k,{c_{out}}}\times\textbf{P}_{\textbf{BS}_{k}})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\mathbb{K}_{i,j,{c_{out}}}\times\textbf{P}_{\textbf{BS}_{i,j}})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\mathbb{K}_{i,j,{c_{out}}}\times(\textbf{P}_{\textbf{B}_{1,j}}+\textbf{P}_{\textbf{S}_{i,j}}))\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\mathbb{K}_{i,j,{c_{out}}}\times(\mathbb{W}_{B}\times\mathbb{P}_{B_{1,j}}\\ &+\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times\mathbb{P}_{S_{m,j}}))\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\mathbb{W}_{B}\times\mathbb{K}_{i,j,{c_{out}}}\times\mathbb{P}_{B_{1,j}}\\ &+\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{i,j,{c_{out}}}\times\mathbb{P}_{S_{m,j}}))\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\mathbb{W}_{B}\times\mathbb{K}_{B_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j}\\ &+\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{i,j,{c_{out}}}\times(\mathbb{P}_{m,j}-\mathbb{P}_{B_{1,j}}))\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j})\\ &+\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{i,j,{c_{out}}}\times\mathbb{P}_{m,j}-\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{i,j,{c_{out}}}\times\mathbb{P}_{B_{1,j}}))\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j})\\ &+\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{i,j,{c_{out}}}\times\mathbb{P}_{m,j}-\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{B_{1,j,{c_{out}}}}\times\mathbb{P}_{m,j}))\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j})\\ &+\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times(\mathbb{K}_{i,j,{c_{out}}}-\mathbb{K}_{B_{1,j,{c_{out}}}})\times\mathbb{P}_{m,j})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j})\\ &+\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{S_{i,j,{c_{out}}}}\times\mathbb{P}_{m,j})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j})\\ &+\sum_{m}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{S}_{m,j,{c_{out}}}}\times\mathbb{P}_{m,j})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j}+\textbf{K}_{\textbf{S}_{i,j,{c_{out}}}}\times\mathbb{P}_{i,j})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}((\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}+\textbf{K}_{\textbf{S}_{i,j,{c_{out}}}})\times\mathbb{P}_{i,j})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{BS}_{i,j,{c_{out}}}}\times\mathbb{P}_{i,j})\\ =&\sum_{k}^{K_{h}\times K_{w}\times C_{in}}(\textbf{K}_{\textbf{BS}_{k,{c_{out}}}}\times\mathbb{P}_{k})\end{aligned}$
[Uncaptioned image] Jinming Cao received her Ph.D. with the School of Computer Science and Technology from Shandong University in 2022. She is currently working as a Research Fellow with the National University of Singapore. Her research interests include computer vision and 3D vision.
[Uncaptioned image] Sicheng Shen is currently an undergraduate student at the Faculty of Science, National University of Singapore. His research interest includes computer vision.
[Uncaptioned image] Qiu Zhou received the Master degree with the School of Computer Science and Tenchnology from Shandong University in 2022. She is currently a Ph.D. student with the Beijing University of Posts and Telecommunications. Her research interests include computer vision and 3D vision.
[Uncaptioned image] Yifang Yin received the BE degree from the Department of Computer Science and Technology, Northeastern University, Shenyang, China, in 2011 and the PhD degree from the National University of Singapore, Singapore, in 2016. She is currently a senior scientist with Institute for Infocomm Research (I2R) in the Agency for Science, Technology and Research, Singapore (A*STAR). She also holds an adjunct faculty position with IIIT-Delhi. Before joining A*STAR, she worked as a senior research fellow with the Grab-NUS AI Lab with the National University of Singapore. She also worked as a Research Intern with the Incubation Center, Research and Technology Group, Fuji Xerox Co., Ltd., Japan, from October, 2014 to March, 2015. Her research interests include machine learning, spatiotemporal data mining, and multimodal analysis in multimedia.
[Uncaptioned image] Yangyan Li received his Ph.D. from University of Chinese Academy of Sciences in 2013. He is currently working in Alibaba group. Before that, he was a professor in Shandong University, and Postdoc in Stanford University and Tel Aviv University. His research interests include computer graphics and computer vision, with an emphasis on 3D learning and vision.
[Uncaptioned image] Roger Zimmermann received the M.S. and Ph.D. degrees from the Viterbi School of Engineering at the University of Southern California. He is a Full Professor at the School of Computing at the National University of Singapore (NUS). He is a Co-PI with the Grab-NUS AI Lab. His research interests are streaming media architectures and protocols for video (DASH) and immersive media (e.g., point cloud data), applications of machine/deep learning, spatiotemporal data management, analytics and applications, and location-based services.

### v2

@@ 4.4. Ablation Study
The two components we propose, ShapeConv and Shape-Architecture, are independent of each other. This means that adding either one of them to the baseline method alone can enhance the model’s performance. To validate the effectiveness of both components, we conducted detailed ablation studies. The results of these studies are presented in Table 4. One can observe that the removal of both ShapeConv and Shape-Architecture leads to degraded model performance. Introducing either the ShapeConv or the Shape-Architecture yields a positive outcome. Combining these two together, i.e., our final ShapeMoiré, achieves the best results, highlighting their complementary effect on the overall model performance.
Table 4. Ablation study of Shapremoiré on the UHDM dataset.
@@
Additionally, we computed the inference time of both the baseline and our ShapeMoiré method. Specifically, we conducted this experiment on a single NVIDIA A40 GPU with 48G memory and ran both models for three runs. Subsequently, we calculated the average inference time for 5,000 images and presented the milliseconds (ms) per image in Table 5. It is evident that our ShapeMoiré method imposes a very negligible overhead on the baseline model in most cases and even runs faster than the baseline due to clock time oscillation. This results validate the efficiency of the proposed ShapeMoiré method.
Table 5. Inference time comparison of baseline and our ShapeMoiré on the UHDM dataset.
@@
(Sun et al., 2018) ShapeMoiré 1.95 2.07 2.21
We conducted experiments with three different settings of the loss weights ($\alpha,\beta$), as shown in Table 6. The results demonstrate that across all configurations, our method consistently outperforms the baseline, indicating that the proposed framework is not highly sensitive to the specific choice of weight parameters and exhibits robust performance. While there are slight performance variations depending on the weighting (e.g., slightly better LPIPS with $\alpha=10$), we attribute these differences to potential dataset-specific characteristics. To maintain fairness and reproducibility, we adopt the balanced setting ($\alpha=1,\beta=1$) for all experiments reported in the main paper.
Table 6. Ablation study on the weighting coefficients for the loss terms. Experiments are conducted using DMCNN as the baseline method on the LCDMoiré dataset.
[…]
Baseline 31.425 .9552 0.093
0.1 1.0 31.975 .9616 0.085
1.0 1.0 31.789 .9562 0.088
10 1.0 32.077 .9609 0.070
4.5. Visualization
Figure 4. Visualization results from the UHDM dataset. Subfigures (a), (b), (c), and (d) respectively showcase different sample cases. In each subfigure, the $I_{moir\acute{e}}$, $I_{gt}$, $Baseline$, and ShapeMoiré correspond to the moiré-contaminated input image, ground truth, baseline output, and ShapeMoiré output, respectively. For each case, $Baseline_{A}$ and ShapeMoiréA denote results based on WDNet (Liu et al., 2020), while $Baseline_{B}$ and ShapeMoiréB are derived using ESDNet-L (Yu et al., 2022) as the baseline. The larger boxes display magnified versions of the corresponding colored boxes, providing a clearer view of fine-grained differences.
In Fig. 4, we present qualitative comparisons across four sample cases—(a), (b), (c), and (d)—each visualized with moiré-contaminated input, ground truth, baseline output, and our ShapeMoiré output. We include two representative baselines, WDNet(Liu et al., 2020) and ESDNet-L (Yu et al., 2022), to assess the generalizability of our method. For each baseline, its corresponding ShapeMoiré variant is denoted as ShapeMoiréA and ShapeMoiréB, respectively. First, we observe that stronger baselines (e.g., ESDNet-L) generally produce better restoration results, as seen by comparing the green boxes of (b) and (d) with those of (a) and (c). However, regardless of the baseline strength, our ShapeMoiré consistently demonstrates superior moiré removal and better structural detail recovery. For instance, in subfigure (a), ShapeMoiréA (blue box) more accurately reconstructs the sailboat and the surrounding sky compared to the baseline (green box). Similarly, in subfigure (b), ShapeMoiréB produces more uniform color on the lake surface. In subfigures (c) and (d), our method further reduces residual moiré artifacts while better preserving textures such as road surfaces. These observations collectively highlight the robustness and effectiveness of our method across varying baseline strengths and image characteristics.
Figure 5. Qualitative results of image demoiréing for real-world cases. (a) Given the same ground truth image $I_{gt}$, different moiré-contaminated inputs $I^{A}_{moir\acute{e}}$ and $I^{B}_{moir\acute{e}}$ are obtained by photographing with different smartphones. (b) and (c) show the demoiréing results for $I^{A}_{moir\acute{e}}$ and $I^{B}_{moir\acute{e}}$ respectively, comparing a baseline method and our proposed ShapeMoiré. Larger boxes represent enlarged views of the corresponding colored small boxes at the borders, providing a clearer visualization of the moiré removal effectiveness.
@@ 4.5. Visualization
To further validate the practicality of our proposed method, we captured real-world images using smartphones and applied both the baseline and ShapeMoiré methods to these cases, as depicted in Fig. 5. It is important to note that the patterns of moiré are not fixed; for the same original image, different devices may capture different moiré patterns, as illustrated by the two images with moiré, where $I^{A}_{moir\acute{e}}$ is the original input image in the UHDM dataset (Yu et al., 2022) and $I^{B}_{moir\acute{e}}$ is from our captured smartphone input image. It can be observed that $I^{A}_{moir\acute{e}}$ exhibits more widespread moiré artifacts across the entire region ($Patttern^{A}$), while in $I^{B}_{moir\acute{e}}$, the artifacts are concentrated in the sky area at the center of the image ($Patttern^{B}$). From the results in Fig. 5, several observations can be made: 1) Our ShapeMoiré consistently outperforms the baseline method, regardless of the two morphologies of moiré, $I^{A}_{moir\acute{e}}$ and $I^{B}_{moir\acute{e}}$; our method exhibits superior moiré removal effects. 2) Our method demonstrates robustness against different moiré patterns, even helping alleviate color distortions. As shown in example $I^{B}_{moir\acute{e}}$ of Fig. 5, due to severe moiré in the blank sky area at the center, noticeable color deviations occur. ShapeMoiré not only excels in removing Moiré patterns but also effectively restores the true colors of the sky.
Figure 6. Qualitative ablation results on the UHDM dataset using ESDNet as the baseline. From left to right: moiré-contaminated input, baseline result, baseline with ShapeConv, ShapeMoiré, and ground truth. The top row of each image shows a zoom-in of the highlighted region, allowing for closer inspection of visual differences.
In Fig. 6, we present qualitative comparisons to visually examine the effects of ablating key components, including ShapeConv and the Shape-Architecture, based on the strong ESDNet(Yu et al., 2022) baseline. As shown, while the baseline already delivers competitive results due to its advanced design, further improvements can still be observed when our proposed modules are added. Specifically, adding ShapeConv leads to better restoration of fine-grained details. For instance, in the shadow area beneath the lamp structure, ShapeConv enables more accurate reconstruction compared to the baseline alone. This suggests that the patch-level shape-aware representation contributes to finer spatial precision. Moreover, incorporating the Shape-Architecture introduces a noticeable enhancement in global image appearance. Compared to the baseline’s overall cooler tone, the Shape-Architecture helps correct color bias, producing a warmer tone that more closely aligns with the ground truth. This demonstrates its ability to recover higher-level structural and color features from the image-level shape stream. Together, these qualitative results visually confirm the effectiveness of each component, serving as a strong perceptual complement to our quantitative evaluations.
5. Conclusion
@@
In this paper, we extend our previously accepted ShapeConv to the image demoiréing domain. Unlike existing methods, our proposed ShapeMoiré addresses the channel inconsistency problem of moiré patterns using shape features. The key extension in the method is the introduction of the Shape-Architecture, which models the shape information from the image-level and acts as a complement to the patch-level ShapeConv. The experimental results demonstrate its superiority over existing state-of-the-arts on four public popular datasets. Furthermore, we verify that our method can be painlessly integrated into four strong baselines to improve their performance, while introducing negligible computational cost during inference. Additionally, our ShapeMoiré yields significant promise in real-world cases with irregular moiré patterns, making it potentially suitable for edge device application and deployment. Beyond demoiréing, we have also applied ShapeMoiré to the image deblurring task and observed promising improvements, suggesting its potential generalizability to broader image restoration problems.
Despite the encouraging results, our method has certain limitations. For instance, the current design primarily targets repetitive moiré patterns and may be less effective for more complex or irregular noise types that do not exhibit clear shape structures. Also, while computational overhead during inference is minimal, training with ShapeConv and Shape-Architecture requires additional resources. In future work, we plan to explore adapting and extending the framework to other image restoration tasks such as general denoising, deraining, and compression artifact removal. Furthermore, integrating more advanced frequency-domain analysis and lightweight model designs could improve both effectiveness and efficiency, making the approach more practical for diverse real-world applications.
Acknowledgements.
This work is fully supported by the Advanced Research and Technology Innovation Centre (ARTIC), the National University of Singapore under Grant (project number: ELDT-RP2).
@@ Appendix A Proof for Equation 9
$\mathbb{F}_{c_{out}}=&\sum_{k}^{K_{h}\times K_{w}\times C_{in}}(\mathbb{K}_{k,{c_{out}}}\times\textbf{P}_{\textbf{BS}_{k}})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\mathbb{K}_{i,j,{c_{out}}}\times\textbf{P}_{\textbf{BS}_{i,j}})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\mathbb{K}_{i,j,{c_{out}}}\times(\textbf{P}_{\textbf{B}_{1,j}}+\textbf{P}_{\textbf{S}_{i,j}}))\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\mathbb{K}_{i,j,{c_{out}}}\times(\mathbb{W}_{B}\times\mathbb{P}_{B_{1,j}}+\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times\mathbb{P}_{S_{m,j}}))\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\mathbb{W}_{B}\times\mathbb{K}_{i,j,{c_{out}}}\times\mathbb{P}_{B_{1,j}}+\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{i,j,{c_{out}}}\times\mathbb{P}_{S_{m,j}}))\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\mathbb{W}_{B}\times\mathbb{K}_{B_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j}+\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{i,j,{c_{out}}}\times(\mathbb{P}_{m,j}-\mathbb{P}_{B_{1,j}}))\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j})+\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{i,j,{c_{out}}}\times\mathbb{P}_{m,j}-\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{i,j,{c_{out}}}\times\mathbb{P}_{B_{1,j}}))\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j})+\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{i,j,{c_{out}}}\times\mathbb{P}_{m,j}-\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{B_{1,j,{c_{out}}}}\times\mathbb{P}_{m,j}))\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j})+\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times(\mathbb{K}_{i,j,{c_{out}}}-\mathbb{K}_{B_{1,j,{c_{out}}}})\times\mathbb{P}_{m,j})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j})+\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}\sum_{m}^{K_{h}\times K_{w}}(\mathbb{W}_{S_{m,i,j}}\times\mathbb{K}_{S_{i,j,{c_{out}}}}\times\mathbb{P}_{m,j})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j})+\sum_{m}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{S}_{m,j,{c_{out}}}}\times\mathbb{P}_{m,j})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}\times\mathbb{P}_{i,j}+\textbf{K}_{\textbf{S}_{i,j,{c_{out}}}}\times\mathbb{P}_{i,j})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}((\textbf{K}_{\textbf{B}_{1,j,{c_{out}}}}+\textbf{K}_{\textbf{S}_{i,j,{c_{out}}}})\times\mathbb{P}_{i,j})\\ =&\sum_{i}^{K_{h}\times K_{w}}\sum_{j}^{C_{in}}(\textbf{K}_{\textbf{BS}_{i,j,{c_{out}}}}\times\mathbb{P}_{i,j})\\ =&\sum_{k}^{K_{h}\times K_{w}\times C_{in}}(\textbf{K}_{\textbf{BS}_{k,{c_{out}}}}\times\mathbb{P}_{k})\end{aligned}$